# Towards a theory of model distillation

### v1

@@ 1.1.1 Defining PAC-distillation
Given a source model $f\in{\mathcal{F}}$ and i.i.d. samples $x_{1},\ldots,x_{n}$ from an unknown distribution ${\mathcal{D}}$, PAC-distillation is the problem of finding a model $g\in{\mathcal{G}}$ whose error approximating $f$ over the distribution is small: error_f,D(g) = P_x D[g(x) f(x)] .
@@
With this definition in hand, we now return to the questions raised above. Before describing our general results on fundamental computational and statistical limits for distillation, we focus on distilling neural networks, since these are of contemporary significance.
@@ 1.1.2 New algorithmic tools to distill neural networks
This constitutes the first polynomial-time algorithm for distillation of neural networks into decision trees. The prior best algorithm for this task uses query access to the neural network model and runs in time $\tilde{O}(d^{2})\cdot s^{O(\log\log s)}$ when the input distribution is assumed to be uniform over the binary hypercube [29]. The caveat in our result is that we additionally require the linear representation property. Nevertheless, we verify experimentally that this property does hold in practice. In Section 3.2.2, we implement our algorithm to successfully extract decision trees from deep neural networks that have been trained to learn a decision tree.
@@ 1.1.3 General computational and statistical theory of distillation
In addition to this result, we show lower bounds on the sample complexity of distillation, prove that the natural analogues of VC-dimension from learning theory do not correctly predict the sample complexity of distillation, and identify a number of new open problems on pinning down the sample complexity of distillation.
@@ 3.2 Provably-efficient distillation of networks into decision trees
Refer to caption Intermediate computations ${\mathcal{G}}_{T}$: $1$ $\neg x_{24}$ $x_{24}$$\neg x_{24}\wedge\neg x_{70}$$\neg x_{24}\wedge x_{70}$ $x_{24}\wedge\neg x_{31}$ $x_{24}\wedge x_{31}$$\neg x_{24}\wedge x_{70}\wedge x_{21}$$\neg x_{24}\wedge x_{70}\wedge x_{21}$ $x_{24}\wedge x_{31}\wedge\neg x_{60}$ $x_{24}\wedge x_{31}\wedge x_{60}$
@@
For any $\epsilon,\delta\in(0,1)$, there is an algorithm that $(\epsilon,\delta)$-distills from ${\mathcal{F}}^{\tau\mbox{-}\mathrm{LRH}}_{{\sf NN},\mathrm{size}\mbox{-}s,\mathrm{depth}\mbox{-}r\mbox{ }\mathrm{trees}}$ to ${\mathcal{F}}_{\mathrm{size}\mbox{-}s,\mathrm{depth}\mbox{-}r\mbox{ }\mathrm{trees}}$ and runs in $\mathrm{poly}(d,m,1/\epsilon,s,2^{r},\log(1/\delta),\tau,B)$ time and $\mathrm{poly}(1/\epsilon,s,\log(d/\delta),\log(\tau B))$ samples from ${\mathcal{D}}$, where $B\geq\max_{{\bm{x}}}\|\varphi_{\theta}({\bm{x}})\|$ is an upper bound on the norm of the network representation. Furthermore, if ${\mathcal{D}}$ is uniform over $\{+1,-1\}$, then there is an algorithm that $(\epsilon,\delta)$-distills from ${\mathcal{F}}^{\tau\mbox{-}\mathrm{LRH}}_{{\sf NN},\mathrm{size}\mbox{-}s,\mathrm{depth}\mbox{-}r\mbox{ }\mathrm{trees}}$ to ${\mathcal{F}}_{\mathrm{size}\mbox{-}s,\mathrm{depth}\mbox{-}r\mbox{ }\mathrm{trees}}$ in $\mathrm{poly}(d,m,1/\epsilon,s,\log(1/\delta),\tau,B)$ time.
@@
Let ${\mathcal{S}}$ be a collection of nondegenerate $k$-clauses. Let ${\mathcal{G}}=\{\mathrm{AND}_{S}\mbox{ for all }S\in{\mathcal{S}}\}$. Suppose that $\varphi:\{0,1\}^{d}\to\mathbb{R}^{m}$ approximately satisfies the $\tau$-LRH with respect to ${\mathcal{G}}$ in the sense that for $g\in{\mathcal{G}}$ there is ${\bm{w}}_{g}\in\mathbb{R}^{m}$ with $\|{\bm{w}}_{g}\|\leq\tau$ such that E_x∼Unif{0,1}^d[(w_g φ(x) - g(x))^2] 2^-k-2 . Then —S2^3k+4 τ^2 E_x∼Unif{0,1}^d[φ(x)^2]  .
@@ 4 Computational theory: a web of reductions
size-$s$, depth-$r$ decision trees $k$-juntas neural networks implicitly computing size-$s$, depth-$r$ decision trees, under LRH neural networks implicitly computing $k$-juntas Prop. 3.2ptProp. 4.5ptThm. 3.6(for uniform dist.)
@@ 5 Statistical theory: bounds on sample complexity
  • Distilling a class ${\mathcal{F}}$ into a class ${\mathcal{G}}$ is no harder than learning ${\mathcal{G}}$. Furthermore, there are some cases where ${\mathcal{G}}$ requires a high number of samples to learn, but we can distill from ${\mathcal{F}}$ to ${\mathcal{G}}$ with no samples; see Section 5.1.
  • The low sample complexity of distillation is a general phenomenon – whenever ${\mathcal{F}}$ is perfectly distillable into ${\mathcal{G}}$ (defined below), very few samples are needed to do so; see Section 5.2.
  • On the other hand, there are cases in which agnostic distillation of ${\mathcal{F}}$ into ${\mathcal{G}}$ can have high sample complexity; see Section 5.3.
  • Finally, we pose the problem of determining a combinatorial property that controls the sample complexity of distillation. We prove that a natural candidate for this property fails; see Section 5.4.
Throughout this section, we view $f\in{\mathcal{F}}$ and $g\in{\mathcal{G}}$ simply as functions, ignoring how they are represented (e.g., neural network, decision tree, or truthtable), which only mattered in previous sections due to the computational constraints om the distillation algorithms.
5.1 Distillation is strictly easier than learning
First we observe that whenever we can (agnostically) learn, we can also (agnostically) distill.
Theorem 5.1 (Distilling is easier than learning).
If concept class ${\mathcal{F}}$ is $(\epsilon,\delta)$-learnable with hypothesis class ${\mathcal{G}}$ in $n$ samples, then source class ${\mathcal{F}}$ is $(\epsilon,\delta)$-distillable into target class ${\mathcal{G}}$ in $n$ samples. Similarly, if hypothesis class ${\mathcal{G}}$ is agnostically $(\epsilon,\delta)$-learnable in $n$ samples, then any source class ${\mathcal{F}}$ can be agnostically $(\epsilon,\delta)$-distilled into target class ${\mathcal{G}}$ in $n$ samples.
@@ 5.1 Distillation is strictly easier than learning
Let ${\mathcal{A}}_{\mathrm{learn}}$ be a learning algorithm that takes in samples $S=(x_{1},\ldots,x_{n})\sim{\mathcal{D}}^{\otimes n}$ and their labels $L=(f(x_{1}),\ldots,f(x_{n}))$, and outputs a hypothesis $g={\mathcal{A}}_{\mathrm{learn}}(S,L)\in{\mathcal{G}}$ such that $\mathbb{P}[\mathrm{error}_{{\mathcal{D}}}(g;f)\leq\varepsilon]\geq 1-\delta$. A distillation algorithm ${\mathcal{A}}_{\mathrm{dist}}$ can get the same guarantee by outputting ${\mathcal{A}}_{\mathrm{dist}}(S,f):={\mathcal{A}}_{\mathrm{learn}}(S,(f(x_{1}),\ldots,f(x_{n})))$. The proof in the agnostic case is analogous. ∎
However, the converse is false: in some cases distillation can require strictly fewer samples than learning. In extreme cases, when the concept class is contained in the hypothesis class, distillation is possible with 0 samples, whereas learning might be sample-expensive.
Theorem 5.2 (Distilling can be strictly easier than learning).
Let ${\mathcal{X}}$ be a finite input space of size $|{\mathcal{X}}|=m$, let ${\mathcal{Y}}=\{0,1\}$, and let ${\mathcal{F}}={\mathcal{G}}={\mathcal{Y}}^{{\mathcal{X}}}$ be all possible functions from ${\mathcal{X}}$ to ${\mathcal{Y}}$. Then: The class ${\mathcal{F}}$ is $(\epsilon=0,\delta=0)$-distillable into ${\mathcal{G}}$ using 00 samples. On the other hand, for any $160<\epsilon\leq 1/$ 16 and for $\delta=1/2$, at least $n\geq\Omega(m/\epsilon)$ samples are needed to $(\epsilon,\delta)$-learn ${\mathcal{F}}$ with ${\mathcal{G}}$.
Proof.
The distillation algorithm ${\mathcal{A}}_{\mathrm{dist}}(S,f)$ that outputs $f$ is valid since $f\in{\mathcal{G}}$ because ${\mathcal{F}}={\mathcal{G}}$. It requires no samples and has error $\epsilon=0$ and soundness $\delta=0$ for any distribution over the inputs.
The lower bound on the number of samples need to learn ${\mathcal{F}}$ is well known: it follows from Theorem 3.5 of [62] because the “VC dimension” of ${\mathcal{F}}$ is $m$. For completeness, we repeat the proof here. Let ${\mathcal{X}}=\{x_{1},\ldots,x_{m}\}$, and let ${\mathcal{D}}$ be the distribution over ${\mathcal{X}}$ that assigns probability mass $1-16\epsilon$ to $x_{1}$ and $16\epsilon/(m-1)$ to each of $x_{2},\ldots,x_{m}$. Let ${\mathcal{I}}={\mathcal{X}}\setminus S$ be the set of inputs that do not appear in the algorithm’s multiset of samples $S$. If the learning algorithm ${\mathcal{A}}_{\mathrm{learn}}$ draws $n\leq(m-1)/(64\epsilon)$ samples, then by a Markov bound
$\mathbb{P}[|{\mathcal{I}}|\leq\frac{m-1}{2}]\geq 3/4.$ (5.1)
The labels of the inputs in ${\mathcal{I}}$ are undetermined by the samples: in particular, there is a concept $f\in{\mathcal{F}}$ agreeing with the samples such that, with respect to this concept, the learning algorithm has error at least $(|{\mathcal{I}}|/2)(16\epsilon/(m-1))$ in expectation. Therefore
$\mathbb{P}[\mathrm{error}({\mathcal{A}}_{\mathrm{learn}}(S,L))\geq\frac{2\epsilon|{\mathcal{I}}|}{m-1}]\geq 3/4.$ (5.2)
By a union bound combining (5.1) and (5.2), we have $\mathbb{P}[\mathrm{error}({\mathcal{A}}_{\mathrm{learn}}(S,L))\geq\epsilon]\geq 1/2$. ∎
In this section, we observe that the low sample complexity of distillation observed in Theorem 5.2 is a general phenomenon. We prove that distillation has a very low sample complexity whenever it is possible to perfectly distill, i.e., when it is possible to carry out the distillation to any given level of accuracy. We define perfect distillation below.
@@ 5.2 Whenever perfect distillation is possible, very few samples are needed
We remark that the definition of perfect distillation does not allow the class ${\mathcal{G}}$ to grow as we vary $\epsilon,\delta$. This means that the results in this section do not apply if, e.g., our scenario is distilling neural networks into logical circuits, and we require larger circuits as $\epsilon$ and $\delta$ decrease. Further research is required to understand the sample complexity in this case; see the open problems in Section 5.4.
First, we consider the case of finite input space, where we show that perfect distillation can always be achieved with 0 samples whenever it is possible.
@@
In the case of countable (potentially infinite) input spaces ${\mathcal{X}}$, we similarly show that perfect distillation can be achieved with very few samples whenever it is possible. The number of samples required is nonzero in this case, but our upper bound scales mildly in $\epsilon,\delta$, and does not depend on any extra structure of ${\mathcal{F}}$ and ${\mathcal{G}}$.
@@ 5.3 Agnostic distillation may nevertheless require a high number of samples
So far, we have shown that distillation is sometimes far easier than learning. But are there situations where distillation still has a high sample complexity? The answer is yes, and we prove it below.
First, we observe that it might not be possible to distill source class ${\mathcal{F}}$ with hypothesis class ${\mathcal{G}}$, even with unboundedly many samples. This is because the functions in ${\mathcal{F}}$ and ${\mathcal{G}}$ may be highly mismatched.
Observation 5.7 (Distillation may be impossible, even with unbounded samples).
Proof.
Agnostic distillation bypasses this obstacle, since here the goal is instead to find a hypothesis in ${\mathcal{G}}$ that approximates the best possible one in ${\mathcal{G}}$. For finite input spaces, agnostic distillation is always possible with a number of samples that grows as fast as the size of the input space ${\mathcal{X}}$, because it can be reduced to agnostic learning by Theorem 5.1. It turns out that in some cases this large number of samples is needed in order to agnostically distill:
@@
Note: this proof shares some elements with the proof that some classes require a large number of samples to agnostically learn (Lemma 3.23 of [73]). This is no coincidence, since lower bounds on the number of samples needed for agnostic distillation also imply lower bounds for agnostic learnability (cf. Theorem 5.1).
5.4 Open problem: combinatorial characterization of sample complexity
We have established that distilling a source class ${\mathcal{F}}$ into a target class ${\mathcal{G}}$ is no harder than the corresponding learning problem, but it may still require a large number of samples or may be impossible. In this section we ask: can we identify a combinatorial property of ${\mathcal{F}}$ and ${\mathcal{G}}$ that characterizes how many samples are needed to distill? We leave this problem open, and provide partial progress by showing that a natural candidate based on the VC dimension fails.
5.4.1 Sample complexity of learning is determined by the VC dimension

### v2

@@ 1.1.1 Defining PAC-distillation
Given a source model $f\in{\mathcal{F}}$ and i.i.d. samples $x_{1},\ldots,x_{n}$ from an unknown distribution ${\mathcal{D}}$, PAC-distillation is the problem of finding a model $g\in{\mathcal{G}}$ whose error approximating $f$ over the distribution is small: $\mathrm{error}_{f,{\mathcal{D}}}(g)=\mathbb{P}_{x\sim{\mathcal{D}}}[g(x)\neq f(x)]\,.$
@@
Remark 1.3 (Relation to the KLCL model of [17]).
The closest setting studied in the literature is the Known-Labeling-Classifier-Learning (KLCL) model of [17]. The KLCL model corresponds to PAC-distillation, but in the special case where ${\mathcal{F}}$ is the set of all possible labeling functions ${\mathcal{X}}\to{\mathcal{Y}}$. In other words, in the KLCL model the labels are known, but they are not required to be instantiated by a classifier in a small class ${\mathcal{F}}$. We discuss this highly-relevant setting further in Section 1.2.
With the definition of PAC-distillation in hand, we now return to the questions raised above. Before describing our general results on fundamental computational and statistical limits for distillation, we focus on distilling neural networks, since these are of contemporary significance.
@@ 1.1.2 New algorithmic tools to distill neural networks
This constitutes the first polynomial-time algorithm for distillation of neural networks into decision trees. The prior best algorithm for this task uses query access to the neural network model and runs in time $\tilde{O}(d^{2})\cdot s^{O(\log\log s)}$ when the input distribution is assumed to be uniform over the binary hypercube [31]. The caveat in our result is that we additionally require the linear representation property. Nevertheless, we verify experimentally that this property can hold in practice. In Section 3.2.2, we implement our algorithm to successfully extract decision trees from deep neural networks that have been trained to learn a decision tree.
@@ 1.1.3 General computational and statistical theory of distillation
Finally, we consider agnostic distillation, where the goal is to distill to a model that competes with the best model in ${\mathcal{G}}$. We provide lower bounds and upper bounds on the sample complexity. Fan Chen and Sasha Rakhlin have pointed out in private communication [38] that the sample complexity of agnostic distillation is equivalent to the sample complexity in the “General Learning Setting” of [99], and therefore it does not admit a characterization via a VC-dimension-like quantity [28].
@@ 1.2 Related work
As previously discussed, the most related work is [17], which formulates the Known-Labeling-Classifier-Learning (KLCL) model. Under this model, the labels are known, but the distribution over the inputs is not. The goal is to find a model that competes with the best possible model in a class ${\mathcal{G}}$. This corresponds to agnostic PAC-distillation, where the source class ${\mathcal{F}}$ is the set of all possible labeling functions. [17] proves a trichotomy in terms of the dependence on $\epsilon$: either 00 samples, $O(1/\epsilon)$, or $\Omega(1/\epsilon^{2})$ samples are needed, and these three regimes correspond to combinatorial properties of the class ${\mathcal{G}}$. Furthermore, [17] proves that there are classes for which learning in the KLCL model is impossible with any finite number of samples. This implies that there are settings in which agnostic PAC-distillation is impossible with any finite number of samples.
@@ 3.2 Provably-efficient distillation of networks into decision trees
Intermediate computations ${\mathcal{G}}_{T}$:
$1$
$\neg x_{24}$
$x_{24}$
$\neg x_{24}\wedge\neg x_{70}$
$\neg x_{24}\wedge x_{70}$
$x_{24}\wedge\neg x_{31}$
$x_{24}\wedge x_{31}$
$\neg x_{24}\wedge x_{70}\wedge x_{21}$
$\neg x_{24}\wedge x_{70}\wedge x_{21}$
$x_{24}\wedge x_{31}\wedge\neg x_{60}$
$x_{24}\wedge x_{31}\wedge x_{60}$
@@
For any $\epsilon,\delta\in(0,1)$, there is an algorithm that $(\epsilon,\delta)$-distills from ${\mathcal{F}}^{\tau\mbox{-}\mathrm{LRH}}_{{\sf NN},\mathrm{size}\mbox{-}s,\mathrm{depth}\mbox{-}r\mbox{ }\mathrm{trees}}$ to ${\mathcal{F}}_{\mathrm{size}\mbox{-}s,\mathrm{depth}\mbox{-}r\mbox{ }\mathrm{trees}}$ and runs in $\mathrm{poly}(d,m,1/\epsilon,s,2^{r},\log(1/\delta),\tau,B)$ time and $\mathrm{poly}(1/\epsilon,s,\log(d/\delta),\log(\tau B))$ samples from ${\mathcal{D}}$, where $B\geq\max_{{\bm{x}}}\|\varphi_{\theta}({\bm{x}})\|$ is an upper bound on the norm of the network representation.
Furthermore, if ${\mathcal{D}}$ is uniform over $\{+1,-1\}$, then there is an algorithm that $(\epsilon,\delta)$-distills from ${\mathcal{F}}^{\tau\mbox{-}\mathrm{LRH}}_{{\sf NN},\mathrm{size}\mbox{-}s,\mathrm{depth}\mbox{-}r\mbox{ }\mathrm{trees}}$ to ${\mathcal{F}}_{\mathrm{size}\mbox{-}s,\mathrm{depth}\mbox{-}r\mbox{ }\mathrm{trees}}$ in $\mathrm{poly}(d,m,1/\epsilon,s,\log(1/\delta),\tau,B)$ time.
@@
Let ${\mathcal{S}}$ be a collection of nondegenerate $k$-clauses. Let ${\mathcal{G}}=\{\mathrm{AND}_{S}\mbox{ for all }S\in{\mathcal{S}}\}$. Suppose that $\varphi:\{0,1\}^{d}\to\mathbb{R}^{m}$ approximately satisfies the $\tau$-LRH with respect to ${\mathcal{G}}$ in the sense that for $g\in{\mathcal{G}}$ there is ${\bm{w}}_{g}\in\mathbb{R}^{m}$ with $\|{\bm{w}}_{g}\|\leq\tau$ such that $\mathbb{E}_{{\bm{x}}\sim\mathrm{Unif}\{0,1\}^{d}}[({\bm{w}}_{g}\cdot\varphi({\bm{x}})-g({\bm{x}}))^{2}]\leq 2^{-k-2}\,.$ Then $|{\mathcal{S}}|\leq 2^{3k+4}\tau^{2}\mathbb{E}_{{\bm{x}}\sim\mathrm{Unif}\{0,1\}^{d}}[\|\varphi({\bm{x}})\|^{2}]\,.$
@@ 4 Computational theory: a web of reductions
size-$s$, depth-$r$ decision trees $k$-juntas neural networks implicitly computing size-$s$, depth-$r$ decision trees, under LRH neural networks implicitly computing $k$-juntas Prop. 3.2Prop. 4.5Thm. 3.6(for uniform dist.)
@@ 5 Statistical theory: bounds on sample complexity
  • Distillation has a low sample complexity – whenever ${\mathcal{F}}$ is perfectly distillable into ${\mathcal{G}}$ (defined below), very few samples are needed to do so. In particular, distilling typically requires fewer samples than learning; see Section 5.1.
  • Nevertheless, agnostic distillation may have a high sample complexity. This follows from results of [17] in the KLCL model of learning, and we give a new proof here; see Section 5.2.
  • Next, we give an upper bound on the sample complexity of agnostic distillation, based on the Pareto Frontier of the classes, which shows that distilling can be much easier than learning in some settings; see Section 5.3.
  • Finally, in personal communication Fan Chen and Sasha Rakhlin [38] have kindly pointed out an equivalence between agnostic distillation and the General Learning Setting of Vapnik [99]. This implies that there is no simple analogue of the VC-dimension that characterizes the sample complexity of agnostic distillation; see Section 5.4.
Throughout this section, we view $f\in{\mathcal{F}}$ and $g\in{\mathcal{G}}$ simply as functions, ignoring how they are represented (e.g., neural network, decision tree, or truthtable), which only mattered in previous sections due to the computational constraints on the distillation algorithms.
In this section, we prove that distillation has a very low sample complexity, whenever it is possible to distill. The qualifier is important, since in some cases distillation might not be possible, even with unboundedly many samples.
Observation 5.1 (Distillation may be impossible, even with unbounded samples).
@@ 5.1 Whenever perfect distillation is possible, very few samples are needed
Therefore, we restrict our attention to the cases in which it is possible to perfectly distill, i.e., when it is possible to carry out distillation to any given level of accuracy. We define perfect distillation below.
@@
We remark that the definition of perfect distillation does not allow the class ${\mathcal{G}}$ to grow as we vary $\epsilon,\delta$. This means that the results in this section do not apply if, e.g., our scenario is distilling neural networks into logical circuits, and we require larger circuits as $\epsilon$ and $\delta$ decrease. Further research is required to understand the sample complexity in this case.
If perfect distillation is possible, we prove that distillation requires very few samples. In the case of finite input spaces, distillation can be achieved with 0 samples.
@@
In the case of countable (potentially infinite) input spaces ${\mathcal{X}}$, the number of samples required might be nonzero. Nevertheless, distillation still requires only $O(\log(1/\delta)/\epsilon)$ samples, which scales mildly in $\epsilon,\delta$, and does not depend on any extra structure of ${\mathcal{F}}$ and ${\mathcal{G}}$.
@@ 5.2 Agnostic distillation may nevertheless require a high number of samples
In this section, we consider the sample complexity of agnostic distillation. Since the goal is to find a hypothesis in ${\mathcal{G}}$ that approximates the best possible one in ${\mathcal{G}}$, we do not restrict to when perfect distillation is be possible.
In contrast with the non-agnostic setting, there are classes ${\mathcal{F}}$ and ${\mathcal{G}}$ such that it is impossible to agnostically distill from ${\mathcal{F}}$ to ${\mathcal{G}}$ using any finite number of samples. This follows from a result in [17, Section 6], which proves the impossibility of agnostically learning certain function classes in the Known-Labeling-Classifier-Learning (KLCL) model. Since the KLCL model is a special case of distillation (where ${\mathcal{F}}$ is the set of all possible functions on ${\mathcal{X}}$), this hardness result extends to PAC-distillation.
Below we give an alternative proof that agnostic distillation may require a high number of samples, focusing on finite input spaces.
@@
Remark 5.7.
To interpret this result, observe that agnostic distillation from $\mathcal{F}$ to $\mathcal{G}$ is no harder than agnostic learning of $\mathcal{G}$. Therefore, it is always possible with a number of samples that grows linearly with the input space size $|{\mathcal{X}}|$. The above theorem shows that in some cases this large number of samples is needed in order to agnostically distill.
@@
This proof shares some elements with the proof that some classes require a large number of samples to agnostically learn (Lemma 3.23 of [76]). This is no coincidence, since lower bounds on the number of samples needed for agnostic distillation also imply lower bounds for agnostic learnability.
5.3 Bounding sample complexity of agnostic distillation using Pareto frontier
Since the sample complexity of agnostic distillation may be large in some cases, and small in others, it is natural to ask whether there is a simple combinatorial quantity that controls the sample complexity.